OpenStudy (loser66):

How to solve? \(x"'+x'(x^2-2)=0\) Please, help

OpenStudy (daniel.ohearn1):

Is that a third derivative? x''' ?

OpenStudy (loser66):

Yes,

OpenStudy (loser66):

Thanks anyway, I think I got it.

OpenStudy (daniel.ohearn1):

What did you get?

OpenStudy (loser66):

Actually, it comes from traveling wave equation, which is u(x,t) = f(x-ct) find the solution for \(u_t+u^2u_x+u_{xxx}=0\)

OpenStudy (loser66):

This is what I have \(u_t= -cf'\\u_x=f'\\u_{xxx}=f"'\) Combine all, I have \(f"' =cf' -f^2f'\) Integral both sides I have

OpenStudy (loser66):

\(f" = cf -\dfrac{f^3}{3}\)

OpenStudy (loser66):

multiple both sides by f', then take integral both sides again, I have \(\dfrac{f'^2}{2}=\dfrac{cf^2}{2}-\dfrac{f^4}{12}\)

OpenStudy (loser66):

Simplify and factor f^2 out, transfer to the LHS \((\dfrac{f'}{f})^2=c-f^2/6\)

OpenStudy (loser66):

integral both sides again, \(ln f = cf -f^3/18\)

OpenStudy (loser66):

so, from here, just solve cubic polynomial.